The everyshi package^{∗}

Martin Schröder

Friedrich-Humbert-Straße 124 D-28759 Bremen MS@Dream.HB.North.DE (INTERNET)

1995/01/25

Abstract

This package defines a new command \EveryShipout analogous to \AtBeginDocument etc., whose argument is executed before each \shipout.

Contents

1 Introduction 1

2 Usage 1 3 Options 1

4 Required packages 2

1 Introduction

This package provides the hook \EveryShipout whose argument is executed after the output routine has constructed \box255, and before \shipout is called.

An example application for this package would be a package for adding text to the bottom of each page. Such a package does exist: prelim2e.

2 Usage

\EveryShipout{code} declares ⟨code⟩ that is saved internally and executed before

each \shipout.

The ⟨code⟩ is executed after \box255 has been constructed by the output routine and can change \box255. \shipout is called after ⟨code⟩.

Repeated use of this command is permitted: the code in the argument is stored (and executed) in the order of their declarations.

3 Options

The package has no options.

∗The version umber of this file is v2.00, last revised 1995/01/25.

The name everyshi is a tribute to the 8+3 file-naming convention of certain “operating systems”; strictly speaking it should be everyshipout.

4 REQUIRED PACKAGES 2

4 Required packages

The package does not require any further packages.